You have issued a total of 450 shares of common stock. If you were given the following account balances, explain how you would figure out (1) the par value per share of stock and (2) the average amount you received per share when you issued the stock?

            Common Stock                                       900

            APIC: Issuance of Common Stock      8,325
Answer #1

Common stock will be recorded at par value at the time of issue any excess amount received will go to additional paid in capital (APIC)

Requirement (1)

Par value of stock = 900/450 shares = $2

Requirement (2)

The average amount received per share at time of issue =

(900+8325)/450 shares = $20.50

Note : please comment down for any further clarifications